A THERMOMETER HAS WRONG CALIBRATION,. IT READS THE MELTING POINT OF ICE -10 DEGREE. IT READS 60 DEGREE IN PLACE OF 50 DEGREE CALCULATE THE TEMPERATURE OF BOILING POINT OF WATER ON THIS SCALE. I KNOW THAT IF THERE IS SUCH A QUESTION THE FORMULA IS C/100= 2 -1/N WHERE 1 IS THE LOWER FIXED POINT ON FAULTY THERMOMETER AND 2 READING ON THE FAULTY THERMOMETER, N= NO OF DIVISIONS BETWEEN UPPER AND LOWER POINTS
 MORE I DON'T KNOW HOW TO WORK OUT PLEASE HELP.


Grade:10

1 Answers

Aniket Rout
6 years ago
Answer should be 130°C.(60-(-10))/(B.P -(-10))= (50-0)/(100-0)So, B.P =130°C. Hope u have understood the answer.........................